How many grams of NH3 is produced if you react 42g N2?
1
Expert's answer
2020-11-17T10:35:12-0500

Equation of reaction

3H2(g)+N2(g)=2NH3(g)

Moles of N2(g) that reacted

Moles=mass/RFM

=42g/28.02g/mole

=1.4958moles

Mole Ratio

N2 : NH3

1 : 2

Moles of NH3= 2*1.4989= 2.9978moles

RMM of NH3= 17.03g/mole

Mass of NH3

Mass = moles*RMM

=17.03g/mole*2.9978moles

=51.05g of NH3(g)